If the edges of a 4-inch by 6-inch photograph are each enlarged by 20%, what is the perimeter of the new photograph? Please, could you help me break down this problem step by step?
Answer by ikleyn(52778)

.
New dimensions are:

    The length = 6*1.2 = 7.2 inches.

    The width = 4*1.2 = 4.8 inches.

Now add 7.2 + 4.8 + 7.2 + 4.8 to get your answer.